OpenStudy (anonymous):

differentiate the following: ((cos(x)(e^x-e^a))/(x-a)(e^x))-((sin(x)*ln(x-a)(e^x-e^a))/(e^x(x-a))

OpenStudy (dumbcow):

haha ... use wolfram :)

OpenStudy (dumbcow):

also you need to differentiate numerator and denominator separately

OpenStudy (anonymous):

I did use Wolfram and boy, it disagreed with both you and I. I'm very, very lost now.

OpenStudy (anonymous):

No, I hadn't gotten that far yet.

OpenStudy (anonymous):

I'd moved on to another impossible problem. Anyway, backtracking now.

OpenStudy (anonymous):

so what I saw on Wolfram had no denominator, that because you're differentiating the denominator from the numerator seperately and you just sent me the numerator?

OpenStudy (dumbcow):

yes correct...the denominator should be easier

OpenStudy (anonymous):

wow, this just reeks of IMPOSSIBLE

OpenStudy (anonymous):

how do I take that monstrosity and get an appreciable answer?

OpenStudy (dumbcow):

the result in Wolfram? plug in x=a and see what it simplifies to ..should get a bunch of zeros

OpenStudy (anonymous):

yeah, well okay, I can see there's a lot of x-a's. So my numerator is essentially e^xcos(x) would you say?

OpenStudy (dumbcow):

correct ... and the denominator is ?

OpenStudy (anonymous):

hmmm, it looks like e^x(x-a)

OpenStudy (anonymous):

right? (u)e^x *(1), u = (x-a)?

OpenStudy (anonymous):

f'(x) of e^x is e^x, right?

OpenStudy (anonymous):

so much to remember algebraically I forget all the smooth and easy stuff

OpenStudy (dumbcow):

yep ...du = 1

OpenStudy (anonymous):

great, so I'm left with (e^x (cos x))/e^x(x-a)

OpenStudy (anonymous):

the e^x's = 1, so I'm finished with cos x/(x-a)?

OpenStudy (anonymous):

x-a = 0, so I have to differentiate a 3rd time?

OpenStudy (anonymous):

cos X/0

OpenStudy (dumbcow):

not quite ... product rule says you add --> e^x + e^x(x-a)

OpenStudy (anonymous):

oh, it's so easy to forget the simple stuff after fighting with all these SYMBOLS for so long. What have us humans come to?

OpenStudy (dumbcow):

haha :)

OpenStudy (dumbcow):

so after applying limit ...x=a --> e^a cos(a) / e^a = cos(a)

OpenStudy (anonymous):

whoa! Not what I am seeing. I have (before replacing x) (e^x(Cosx))/(2e^x(x-a))

OpenStudy (anonymous):

maybe I did that wrong, maybe the denominator is (e^x(x-a))+e^x?

OpenStudy (dumbcow):

yes thats it....you can't combine them to make 2e^x

OpenStudy (anonymous):

Perrrrrrfect! You are uber patient.

OpenStudy (anonymous):

Cos a is my answer then!

OpenStudy (dumbcow):

alright good you see it now ...yep